# Conformal prediction with local weights: randomization enables robust guarantees

### v4

@@ 6.2.1 Derandomizing RLCP via $m$-randomly-localized CP ($m$-RLCP)
We will refer to this method as $m$-randomly-localized conformal prediction ($m$-RLCP).
@@
To summarize, this intuitive approach towards derandomizing RLCP shows substantial drawbacks in terms of empirical performance (as well as weaker theory as compared to RLCP). Thus, finding a better way to aggregate across multiple runs of RLCP, to achieve a less random method without compromising on theoretical properties or empirical performance, remains an important open question.
@@ A.3 Asymptotic test-conditional coverage
In particular, if we assume $h_{n}\to 0$, and $h_{n}\cdot n^{1/d}\to\infty$, and if we also assume that $f_{X}(x)$ is bounded away from zero in a neighborhood of $x=x_{0}$, and that that $x\mapsto P_{S|X}(\cdot|x)$ is continuous at $x=x_{0}$ with respect to the total variation distance, meaning that
@@ C.3.1 Local coverage experiments with additional bandwidths
In Section 6.1, we have seen how localization based approaches, such as RLCP, and calibration based approaches, such as the method of Gibbs et al., (2023), fit into a more general framework of conformal algorithms. As mentioned before, the localization step of RLCP with a particular bandwidth, (Step (b) of the three-step workflow described in Section 6.1), and the calibration step of Gibbs et al., (2023) with respect to a particular functional class $\mathcal{F}$ (Step (c) of the workflow) can be viewed as approximately interchangeable—both approaches serve to approximate test-conditional coverage via a local relaxation. In particular, the two approaches lead to a similar tradeoff between localization and sample size. Choosing a richer class of functions $\mathcal{F}$ (for the method of Gibbs et al., (2023)) or a smaller bandwidth for the kernel $H$ (for RLCP) would have similar effects on the outcome, by placing a stronger requirement on the coverage properties of the algorithm and requiring more adaptivity to local properties of the data; on the other hand, for both types of methods, this would come at a cost of reduced effective sample size, and therefore, the potential to return wider or less informative prediction intervals.
In order to compare these methods, and better understand how they share the same tradeoff, we repeat the univariate experiment of Section 5.1.1 (Setting 1). [^9] The method of Gibbs et al., (2023) is run using the code provided in that paper, which is available at https://github.com/jjcherian/conditional-conformal. In Figure C.12 we plot the local coverage (as defined in Section 5.1.1) for each method, averaged over $50$ independent trials. Both Gibbs et al., (2023)’s method and RLCP are implemented in two versions, “high adaptivity” and “low adaptivity”. For Gibbs et al., (2023)’s method, the function class is chosen to be more fine or more coarse,

### v5

@@ 6.2.1 Derandomizing RLCP via $m$-RLCP
We will refer to this method as $m$-RLCP.
@@
To summarize, this intuitive approach towards derandomizing RLCP shows substantial drawbacks in terms of empirical performance (as well as weaker theory as compared with RLCP). Thus, finding a better way to aggregate across multiple runs of RLCP, to achieve a less random method without compromising on theoretical properties or empirical performance, remains an important open question.
@@ A.3 Asymptotic test-conditional coverage
In particular, if we assume $h_{n}\to 0$, and $h_{n}\cdot n^{1/d}\to\infty$, and if we also assume that $f_{X}(x)$ is bounded away from zero in a neighbourhood of $x=x_{0}$, and that that $x\mapsto P_{S|X}(\cdot|x)$ is continuous at $x=x_{0}$ with respect to the total variation distance, meaning that
@@ C.3.1 Local coverage experiments with additional bandwidths
In Section 6.1, we have seen how localization based approaches, such as RLCP, and calibration based approaches, such as the method of Gibbs et al., (2023), fit into a more general framework of conformal algorithms. As mentioned before, the localization step of RLCP with a particular bandwidth, (Step (b) of the three-step workflow described in Section 6.1), and the calibration step of Gibbs et al., (2023) with respect to a particular functional class $\mathcal{F}$ (Step (c) of the workflow) can be viewed as approximately interchangeable—both approaches serve to approximate test-conditional coverage via a local relaxation. In particular, the two approaches lead to a similar trade-off between localization and sample size. Choosing a richer class of functions $\mathcal{F}$ (for the method of Gibbs et al., (2023)) or a smaller bandwidth for the kernel $H$ (for RLCP) would have similar effects on the outcome, by placing a stronger requirement on the coverage properties of the algorithm and requiring more adaptivity to local properties of the data; on the other hand, for both types of methods, this would come at a cost of reduced effective sample size, and therefore, the potential to return wider or less informative prediction intervals.
In order to compare these methods, and better understand how they share the same trade-off, we repeat the univariate experiment of Section 5.1.1 (Setting 1). [^9] The method of Gibbs et al., (2023) is run using the code provided in that paper, which is available at https://github.com/jjcherian/conditional-conformal. In Figure C.12 we plot the local coverage (as defined in Section 5.1.1) for each method, averaged over $50$ independent trials. Both Gibbs et al., (2023)’s method and RLCP are implemented in two versions, “high adaptivity” and “low adaptivity”. For Gibbs et al., (2023)’s method, the function class is chosen to be more fine or more coarse,